# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v2

@@ I Introduction
Safe autonomous robot navigation consists in steering a robot to a target location while avoiding obstacles. One commonly used navigation technique is the Artificial Potential Field (APF) approach [1], where a combination of attractive and repulsive vector fields guides the robot safely to the target location. However, this approach faces challenges with certain obstacle arrangements, leading to undesired local minima. The Navigation Function (NF) approach [2] is effective in sphere world environments [3, 4], addressing the local minima issue by limiting the repulsive field’s influence around the obstacles by means of a properly tuned parameter. This method ensures almost [^1] Almost global convergence here refers to the convergence from all initial conditions except a set of zero Lebesgue measure. global convergence of the robot to the target location. To apply the NF approach to environments with general convex and star-shaped obstacles, diffeomorphic mappings from [3] and [5] can be used. However, these mappings require global knowledge of the environment for implementation. The NF-based approach has been extended in [6] to environments containing convex obstacles with smooth boundaries. The authors established sufficient conditions on the curvature of the obstacles’ boundaries to guarantee almost global convergence to a neighborhood surrounding the a priori unknown target location. However, it is assumed that the shapes of the obstacles avoided by the robot are known.
@@
In [7], the authors proposed a feedback controller based on Nagumo’s theorem [8, Theorem 4.7], for autonomous navigation in environments with general convex obstacles. The forward invariance of the obstacle-free space is ensured by projecting the ideal velocity control vector (pointing to the target) onto the tangent cone at the boundary of the obstacle whenever it points to the obstacle. In [9], a control barrier function-based approach was used for robot navigation in an environment with a single spherical obstacle. In [10], the authors proposed a separating hyperplane-based autonomous navigation algorithm for the robot navigating in environments with convex obstacles with smooth boundaries. This approach was extended in [11] to handle partially known environments with non-convex obstacles, where it is assumed that the robot possesses geometrical information about the non-convex obstacles but lacks knowledge of their precise locations in the workspace. However, due to the topological obstruction to global asymptotic stability, with continuous time-invariant vector fields, in sphere words [2], the above-mentioned approaches provide at best almost global asymptotic stability guarantees.
In [12], hybrid control techniques were employed to achieve global stabilization of the target location in environments with sufficiently separated ellipsoidal obstacles. In [13], the authors utilized the hybrid control techniques to propose an autonomous navigation technique for a nonholonomic robot operating in environments with non-convex obstacles under some restrictions on the inter-obstacle arrangements. In [14], the authors proposed a discontinuous feedback control law for autonomous robot navigation in partially known two-dimensional environments. When a known obstacle is encountered, the control vector aligns with the negative gradient of the navigation function. However, when close to an unknown obstacle, the robot moves along its boundary, relying on the local curvature information of the obstacle. Similar to [13], this approach is applicable to two-dimensional environments and assumes smooth obstacle boundaries without sharp edges. In our earlier work [15], we proposed a hybrid feedback control strategy to address the problem of autonomous robot navigation in two-dimensional environments with arbitrarily-shaped convex obstacles.
@@
In this paper, we propose a hybrid control strategy, endowed with global asymptotic stability guarantees, for autonomous robot navigation in three-dimensional environments with arbitrarily-shaped convex obstacles that can have non-smooth boundaries. We also propose a sensor based version of our approach for autonomous navigation in a priori unknown three-dimensional environments with arbitrarily-shaped convex obstacles.
@@ III Problem Formulation
The minimum separation between any pair of obstacles should be greater than $2r$ i.e., for all $i,j\in\mathbb{I},i\neq j,$ one has
$d(\mathcal{O}_{i},\mathcal{O}_{j}):=\underset{\mathbf{p}\in\mathcal{O}_{i},\mathbf{q}\in\mathcal{O}_{j}}{\text{min}}\|{\mathbf{p}-\mathbf{q}}\|>2r.$ (8)
According to Assumption 8 and the compactness of the obstacles, there exists a minimum separating distance between any pair of obstacles $\bar{r}=\underset{i,j\in\mathbb{I},i\neq j}{\min}d(\mathcal{O}_{i},\mathcal{O}_{j})>2r$. Moreover, for collision-free navigation we require $d(\mathbf{0},\mathcal{O}_{\mathcal{W}})-r>0$. We define a positive real $\bar{r}_{s}$ as
$\bar{r}_{s}=\underset{}{\min}\left\{\frac{\bar{r}}{2}-r,d(\mathbf{0},\mathcal{O}_{\mathcal{W}})-r\right\}.$ (9)
We then pick an arbitrarily small value $r_{s}\in(0,\bar{r}_{s})$ as the minimum distance that the robot should maintain with respect to any obstacle.
The obstacle-free workspace is then defined as $\mathcal{W}_{0}:=\mathcal{W}\setminus\mathcal{O}_{\mathcal{W}}.$ Given $y\geq 0$, a $y-$eroded obstacle-free workspace, $\mathcal{W}_{y}$ is defined as
@@
    Safety: the robot-centered obstacle-free workspace $\mathcal{W}_{r_{a}}$ is forward invariant,
@@
    Global Asymptotic Stability: Any given target location $\mathbf{x}=\mathbf{0}\in\mathcal{W}_{r_{a}}^{\circ}$ is a globally asymptotically stable equilibrium for the closed-loop system.
@@ IV Hybrid Control For Obstacle Avoidance
In the proposed scheme, the robot operates in two modes based on the value of a mode indicator variable $m\in\mathbb{M}:=\{0,1\}$. The move-to-target mode $(m=0)$ is adopted when the robot is away from the obstacles and the the obstacle-avoidance mode $(m=1)$ is adopted when the robot is close to an obstacle obstructing its motion towards the target location. In the move-to-target mode, the robot moves straight towards the target location. In the obstacle-avoidance mode, the robot navigates around the obstacle while staying withing its $\gamma-$neighborhood, where $\gamma\in(0,\bar{r}_{s}-r_{s})$. When the robot operates in the obstacle-avoidance mode, the robot’s center has a unique closest point on the nearest obstacle. Furthermore, to prevent the robot from getting trapped in a loop around an obstacle, the proposed obstacle-avoidance strategy confines the robot’s center to a hyperplane that passes through the target location. This ensures that the robot will eventually reach a position where the nearest obstacle does not intersect with the line segment joining the center of the robot and the target location.
@@ IV-A Hybrid control design
where $\mathbf{I}_{3}\in\mathbb{R}^{3\times 3}$ is the identity matrix and $\mathbf{x}_{\pi}:=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}).$ The operator $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ provides the closest point on the obstacle-occupied workspace from the location of the center of the robot $\mathbf{x}$, as defined in Section II-B. Notice that, since the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$ are convex and the parameter $\gamma\in(0,\bar{r}_{s}-r_{s})$, according to (9), the robot will have a unique closest point to the obstacles whenever its center is in the $(r_{a}+\gamma)-$neighborhood of these obstacles. On the other hand, there may be some locations in the $(r_{a}+\gamma)-$neighborhood of the obstacle $\mathcal{O}_{0}=(\mathcal{W}^{\circ})^{c}$ for which the uniqueness of the closest point from robot’s center to the obstacle $\mathcal{O}_{0}$ cannot be guaranteed. However, as discussed later in Remark 1, the design of the flow sets and jump sets guarantees that the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{a})$ is never activated in the region $\mathcal{N}_{r_{a}+\gamma}(\mathcal{O}_{0}).$
@@
$\mathbf{P}(\mathbf{a}):=\mathbf{I}_{3}-\frac{\mathbf{a}\mathbf{a}^{\intercal}}{\lVert\mathbf{a}\rVert^{2}}.$ (14)
For any vector $\mathbf{x}\in\mathbb{R}^{3},$ the vector $\mathbf{P}(\mathbf{a})\mathbf{x}$ corresponds to the projection of $\mathbf{x}$ onto the hyperplane orthogonal to $\mathbf{a}$. As discussed later in Section IV-C, the coordinates of the unit vector $\mathbf{a}$ are updated when the robot switches from the move-to-target mode to the obstacle-avoidance mode according to the update law $\mathbf{L}(\xi)$, whose design is provided later in Section IV-C.
@@ IV-B Geometric construction of the flow and jump sets
Next, we define an exit region $\mathcal{R}_{e}$ as the part of the $\gamma-$neighborhood of the $r_{a}-$dilated obstacles that do not belong the landing region. The exit region $\mathcal{R}_{e}$ is defined as
@@
Figure 1: The geometrical representation of the landing region and the exit region.
@@ IV-B2 Flow and jump sets (obstacle-avoidance mode)
$\mathcal{J}_{1}^{\mathcal{W}}:=\left(\mathcal{W}\setminus\mathcal{D}_{r_{a}+\gamma}^{\circ}(\mathcal{O}_{\mathcal{W}})\right)\cup\mathcal{ER}^{\mathbf{h}}{\cup\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))}.$ (22)
@@
$\mathcal{F}_{1}:=\{\xi\in\mathcal{K}|\mathbf{x}\in\mathcal{F}_{1}^{\mathcal{W}},m=1,s\neq s_{0}\},$ (25)
@@
where $\mathcal{J}_{\mathbf{x}}:=\{\xi\in\mathcal{K}|m=1,\mathbf{x}\in\mathcal{J}_{1}^{\mathcal{W}}\}$ and $\mathcal{J}_{s}:=\{\xi\in\mathcal{K}|m=1,s=s_{0}\}$ with $s_{0}\in\mathbb{R}_{\geq 0}$ being the initial value of the state $s$, i.e., $s(0,0)=s_{0}$.
@@
IV-C Update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m)$
@@
$\mathbf{L}_{0}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)=\begin{bmatrix}\mathbf{x}\\ \mathbf{A}(\mathbf{x})\\ 1\\ s+1\end{bmatrix}.$ (28)
Given $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$, the set-valued mapping $\mathbf{A}:\mathbb{R}^{3}\rightrightarrows\mathbb{S}^{2}$ is defined as
$\mathbf{A}(\mathbf{x})=\begin{cases}\mathbf{q}\in\mathcal{P}^{\perp}(\mathbf{x}),&\mathbf{x}^{\times}\mathbf{x}_{\pi}=\mathbf{0},\\ \mathbf{q}\in\frac{\mathbf{x}^{\times}\mathbf{x}_{\pi}}{\lVert\mathbf{x}^{\times}\mathbf{x}_{\pi}\rVert},&\mathbf{x}^{\times}\mathbf{x}_{\pi}\neq\mathbf{0},\end{cases}$ (29)
@@
$\mathbf{L}_{1}(\mathbf{x},\mathbf{h},\mathbf{a},1,s)=\begin{bmatrix}\mathbf{h}\\ \mathbf{a}\\ 0\\ s+1\end{bmatrix}.$ (31)
@@
Since the parameter $\gamma\in(r_{s},\bar{r}_{s}-r_{s})$, where $\bar{r}_{s}$ is defined in (9), the $(r_{a}+\gamma)-$dilated obstacles $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, $\forall i\in\mathbb{I}$, are disjoint, and the target location $\mathbf{0}\in\mathcal{W}_{r_{a}+\gamma}^{\circ}$. Furthermore, according to (24), the set $\mathcal{F}_{1}^{\mathcal{W}}$ is contained within the region $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{\mathcal{W}})$. Hence, the proposed control law enables the robot to avoid one obstacle at a time.
@@
  • Parameters selection: the target location is set at the origin with $\mathbf{0}\in\mathcal{W}_{r_{a}}^{\circ}$, and $\xi(0,0)\in\mathcal{K}$. The gain parameters $\kappa_{s}$ and $\kappa_{r}$ are set to positive values, and sufficiently small positive value is chosen for $\bar{\epsilon}$, used in (23). The scalar parameter $\gamma$, used in the construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, is selected such that $\gamma\in(0,\bar{r}_{s}-r_{s})$, where $\bar{r}_{s}$ is defined in (9). The parameters $\gamma_{a}$ and $\gamma_{s}$ are set to satisfy $0<\gamma_{a}<\gamma_{s}<\gamma.$

### v3

@@ I INTRODUCTION
Safe autonomous robot navigation consists in steering a robot to a target location while avoiding obstacles. One common method is the Navigation Function (NF) approach [1], where a combination of attractive and repulsive vector fields, with a properly tuned parameter, guarantees safe, almost [^1] Almost global convergence refers to convergence from all initial locations except a set of zero Lebesgue measure. global convergence of the robot to the target location in sphere worlds. For general convex and star-shaped obstacles, diffeomorphic mappings [1, 2] allow the use of the NF approach but require complete knowledge of the environment. In [3], the NF method has been extended to convex obstacles that satisfy certain curvature properties, and are not too flat and not too close to the target. The work in [4] removes this flatness limitation, in environments with ellipsoidal obstacles, by using the Hessian to locally transform areas near obstacles into spherical regions. However, similar to [3], it is assumed that the entire shape of the obstacle is known when the robot visits its neighbourhood.
In [5, 6, 7], the authors combine control Lyapunov functions (CLF) and control barrier functions (CBF) to design feedback laws that ensure safe convergence to a target location. The authors in [6] introduced a CBF-based method for multi-robot navigation in two-dimensional environments with circular obstacles.A comparative analysis between the APF and CBF approaches can be found in [7]. As shown in [8] and [9], CLF-CBF methods, like the NF approach, suffer from the undesired equilibria problem and guarantee, at best, almost global asymptotic stability in sphere worlds.
@@
In [10], a feedback controller based on Nagumo’s theorem [11, Theorem 4.7] was proposed for autonomous navigation around general obstacles. The method ensures forward invariance of the obstacle-free space by projecting the ideal velocity control vector, which points towards the target, onto the tangent cone at the obstacle boundary when necessary. The research in [10] was extended in [12] to guarantee almost global asymptotic stability in unknown environments with strongly convex obstacles. The work in [13] introduced a reactive navigation approach using separating hyperplanes for robots in environments with unknown, sufficiently separated, strongly convex obstacles. This was further extended in [14] to handle partially known environments with non-convex obstacles, where the robot has information about the shape of the non-convex obstacles but not their locations in the workspace.
However, the approaches discussed above provide, at best, almost global convergence guarantees due the undesired equilibria that are generated when using continuous time-invariant vector fields [15]. This can be resolved by introducing discontinuities in the control, as shown in [16, 17, 18, 19, 20].
In [16], a discontinuous feedback law was developed for robot navigation in partially known two-dimensional environments, guiding the robot with the navigation function gradient near known obstacles and boundary-following near unknown ones. In [17], the authors proposed a discontinuous feedback controller for autonomous navigation of nonholonomic robots in two-dimensional environments with non-convex obstacles, imposing restrictions on the curvatures of obstacles’ boundaries and the inter-obstacle arrangements. In [18], hybrid control methods are employed to achieve robust global asymptotic convergence of a group of planar robots to the maximizer of an unknown potential field, whose value can be measured at any location, while avoiding a single obstacle in a 2-dimensional environment. In [19], the authors proposed a hybrid control law to globally asymptotically stabilize a class of linear systems with drift while avoiding neighbourhoods of unsafe isolated points. In [20], hybrid control techniques were employed to achieve global stabilization of target locations in $n$-dimensional environments with sufficiently separated ellipsoidal obstacles.
@@
In our earlier work [21, 22], we proposed a hybrid feedback control strategy for robot navigation in two-dimensional environments with convex and non-convex obstacles. In this paper, we extend this approach to ensure autonomous robot navigation in unknown $3$-dimensional environments with arbitrarily-shaped convex obstacles, including those with non-smooth boundaries. Our solution guarantees safe, global asymptotic convergence of the robot to the target location, unlike the methods in [4, 12, 13], which guarantee, at best, almost global asymptotic convergence. The approach in [10] applies to environments with ellipsoidal obstacles, the work in [12, 13] is limited to strongly convex obstacles, and in [16, 17] the obstacles are required to have smooth boundaries. In contrast, the proposed hybrid feedback controller works for arbitrarily-shaped convex obstacles with no restrictions on inter-obstacle arrangement, except for a mild separation condition. We also propose a sensor-based version of our approach for autonomous navigation in a priori unknown $3$-dimensional environments with arbitrarily-shaped convex obstacles.
@@ III PROBLEM FORMULATION
The minimum separation between any pair of obstacles should be greater than or equal to $2(r+\delta)$ i.e., for all $i,j\in\mathbb{I},i\neq j,$ one has
$d(\mathcal{O}_{i},\mathcal{O}_{j}):=\underset{\mathbf{p}\in\mathcal{O}_{i},\mathbf{q}\in\mathcal{O}_{j}}{\text{min}}\|{\mathbf{p}-\mathbf{q}}\|\geq 2(r+\delta).$ (8)
We then pick an arbitrary value $r_{s}\in(0,\delta)$ as the minimum distance that the robot should maintain with respect to any obstacle. The obstacle-free workspace is then defined as $\mathcal{W}_{0}:=\mathcal{W}\setminus\mathcal{O}_{\mathcal{W}}.$ Given $y\geq 0$, a $y-$eroded obstacle-free workspace, $\mathcal{W}_{y}$ is defined as
@@
    Safety: the obstacle-free workspace $\mathcal{W}_{r_{a}}$ is forward invariant with respect to the center of the robot,
@@
    Global Asymptotic Stability: Any given target location $\mathbf{x}^{\mathbf{d}}\in(\mathcal{W}_{r_{a}})^{\circ}$ is a globally asymptotically stable equilibrium for the closed-loop system. Without loss of generality, we will consider $\mathbf{x}^{\mathbf{d}}=\mathbf{0}$.
@@ IV HYBRID CONTROL FOR OBSTACLE AVOIDANCE
In the proposed scheme, the robot operates in two modes based on the value of a mode indicator variable $m\in\mathbb{M}:=\{0,1\}$. The move-to-target mode $(m=0)$ is adopted when the robot is away from the obstacles and the obstacle-avoidance mode $(m=1)$ is adopted when the robot is close to an obstacle obstructing its motion towards the target location. In the move-to-target mode, the robot moves straight towards the target location. In the obstacle-avoidance mode, the robot navigates around the obstacle while staying within its $\gamma-$neighborhood, where $\gamma\in(0,\delta-r_{s})$. When the robot operates in the obstacle-avoidance mode, the robot’s center has a unique closest point on the nearest obstacle. Furthermore, to prevent the robot from getting trapped in a loop around an obstacle, the proposed obstacle-avoidance strategy confines the robot’s center to a hyperplane that passes through the target location. It then guides the robot along the boundary to a position where the nearest obstacle no longer intersects with the line segment connecting the robot’s center and the target location, as shown in Fig. 1a.
@@ IV-A Hybrid control design
where $\mathbf{I}_{3}\in\mathbb{R}^{3\times 3}$ is the identity matrix and $\mathbf{x}_{\pi}:=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}}).$ The location $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ is the closest point on the set $\mathcal{O}_{\mathcal{W}}$ from the robot’s center $\mathbf{x}$, as defined in Section II-B. Notice that, since the obstacles $\mathcal{O}_{i},i\in\mathbb{I}\setminus\{0\}$ are convex and the parameter $\gamma\in(0,\delta-r_{s})$, according to Assumption 8, the robot will have a unique closest point to the obstacles whenever its center is in the $(r_{a}+\gamma)-$neighborhood of these obstacles. On the other hand, there may be some locations in the $(r_{a}+\gamma)-$neighborhood of the obstacle $\mathcal{O}_{0}=(\mathcal{W}^{\circ})^{c}$ for which the uniqueness of the closest point from robot’s center to the obstacle $\mathcal{O}_{0}$ cannot be guaranteed. However, as discussed later in Remark 1, the design of the flow sets and jump sets guarantees that the obstacle-avoidance control vector $\mathbf{v}(\mathbf{x},\mathbf{a})$ is never activated in the region $\mathcal{N}_{r_{a}+\gamma}(\mathcal{O}_{0}).$
@@
$\mathbf{P}(\mathbf{a}):=\mathbf{I}_{3}-{\mathbf{a}\mathbf{a}^{\intercal}},$ (13)
where $\mathbf{a}\in\mathbb{S}^{2}$. For any vector $\mathbf{x}\in\mathbb{R}^{3},$ the vector $\mathbf{P}(\mathbf{a})\mathbf{x}$ corresponds to the projection of $\mathbf{x}$ onto the hyperplane orthogonal to $\mathbf{a}$. As discussed later in Section IV-C, the coordinates of the unit vector $\mathbf{a}$ are updated when the robot switches from the move-to-target mode to the obstacle-avoidance mode according to the update law $\mathbf{L}(\xi)$, whose design is provided later in Section IV-C.
@@
Figure 1: (a) Robot trajectory moving along the hyperplane $\mathcal{P}(\mathbf{h},\mathbf{a})$ while avoiding obstacle $\mathcal{O}_{i}$. (b) The geometrical representation of the landing region and the exit region.
@@ IV-B Geometric construction of the flow and jump sets
Next, we define an exit region $\mathcal{R}_{e}$ as the part of the $\gamma-$neighborhood of the $r_{a}-$dilated obstacles that do not belong the landing region, as shown in Fig. 1b. The exit region $\mathcal{R}_{e}$ is defined as
@@ IV-B2 Flow and jump sets (obstacle-avoidance mode)
$\mathcal{J}_{1}^{\mathcal{W}}:=\left(\mathcal{W}\setminus\mathcal{D}_{r_{a}+\gamma}^{\circ}(\mathcal{O}_{\mathcal{W}})\right)\cup\mathcal{ER}^{\mathbf{h}}{\cup\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{0}))}\cup\mathcal{S}_{\mathbf{0}},$ (21)
where the set $\mathcal{S}_{\mathbf{0}}:=\overline{\mathcal{B}_{\gamma}(\mathbf{0})\setminus\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}})}$. The inclusion of the set $\mathcal{S}_{\mathbf{0}}$ in the set $\mathcal{J}_{1}^{\mathcal{W}}$ allows us to ensure the stability of the origin, as stated later in Theorem 1.
@@
$\mathcal{F}_{1}:=\{\xi\in\mathcal{K}|\mathbf{x}\in\mathcal{F}_{1}^{\mathcal{W}},m=1,s\notin(s_{0},s_{0}+\delta_{s})\},$ (24)
@@
where $\mathcal{J}_{\mathbf{x}}:=\{\xi\in\mathcal{K}|m=1,\mathbf{x}\in\mathcal{J}_{1}^{\mathcal{W}}\}$ and $\mathcal{J}_{s}:=\{\xi\in\mathcal{K}|m=1,s=[s_{0},s_{0}+\delta_{s}]\}$ with $s_{0}\in\mathbb{R}_{\geq 0}$ being the initial value of the state $s$, i.e., $s(0,0)=s_{0}$, and $0<\delta_{s}<\tau_{s}$ for some $\tau_{s}>0$.
@@
IV-C Update law $\mathbf{L}(\mathbf{x},\mathbf{h},\mathbf{a},m,s)$
@@
$\mathbf{L}_{0}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)=\left\{\begin{bmatrix}\mathbf{x}\\ \mathbf{a}^{\prime}\\ 1\\ s+\tau_{s}\end{bmatrix},\mathbf{a}^{\prime}\in\mathbf{A}(\mathbf{x})\right\},$ (27)
where $\tau_{s}>0$. Given $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{\mathcal{W}}))$, the set-valued mapping $\mathbf{A}:\mathbb{R}^{3}\rightrightarrows\mathbb{S}^{2}$ is defined as
$\mathbf{A}(\mathbf{x})=\begin{cases}\mathbf{q}\in\mathcal{P}^{\perp}(\mathbf{x}),&\mathbf{x}^{\times}\mathbf{x}_{\pi}=\mathbf{0},\\ \mathbf{q}\in\pm\frac{\mathbf{x}^{\times}\mathbf{x}_{\pi}}{\lVert\mathbf{x}^{\times}\mathbf{x}_{\pi}\rVert},&\mathbf{x}^{\times}\mathbf{x}_{\pi}\neq\mathbf{0},\end{cases}$ (28)
@@
$\mathbf{L}_{1}(\mathbf{x},\mathbf{h},\mathbf{a},1,s)=\begin{bmatrix}\mathbf{h}\\ \mathbf{a}\\ 0\\ s+\tau_{s}\end{bmatrix}.$ (30)
@@
Since the parameter $\gamma\in(r_{s},\delta-r_{s})$, according to Assumption 8, the $(r_{a}+\gamma)-$dilated obstacles $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, $\forall i\in\mathbb{I}$, are disjoint, and the target location $\mathbf{0}\in\mathcal{W}_{r_{a}+\gamma}^{\circ}$. Furthermore, according to (23), the set $\mathcal{F}_{1}^{\mathcal{W}}$ is contained within the region $\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{\mathcal{W}})$. Hence, the proposed control law enables the robot to avoid one obstacle at a time.
@@
  • Parameters selection: the target location is set at the origin with $\mathbf{0}\in\mathcal{W}_{r_{a}}^{\circ}$, and $\xi(0,0)\in\mathcal{K}$. The gain parameters $\kappa_{s}$ and $\kappa_{r}$ are set to positive values, and sufficiently small positive value is chosen for $\bar{\epsilon}$, used in (22). The scalar parameter $\gamma$, used in the construction of the flow set $\mathcal{F}$ and the jump set $\mathcal{J}$, is selected such that $\gamma\in(0,\delta-r_{s})$. The parameters $\gamma_{a}$ and $\gamma_{s}$ are set to satisfy $0<\gamma_{a}<\gamma_{s}<\gamma.$ The parameter $\delta_{s}$, used in (24), is chosen such that $\delta_{s}\in(0,\tau_{s})$, where $\tau_{s}>0$.